determine: (Bonus: +10 pts.) Given: f) = cose + + a. The inverse function, f-(0) b. The graph of the inverse function
Answer #1

Answer fo = caso+ 7/2 a) The inverse function & Co f(0) = Cosotil f(0)_7/2 =coso fo-ah Cost (fo) - 7/4) So o>ftoj f(o) - fo=Lost (n) -y 5 cost (n-02) 2t 2